# Stability of universal properties against perturbations of the Markov Chain Monte Carlo algorithm

### v1

@@ Abstract
We numerically investigate the stability of universal properties at continuous phase transitions against perturbations of the Markov Chain Monte Carlo algorithm used to simulate the system. We consider the three dimensional XY model as test bed, and both local (single site Metropolis) and global (single cluster) updates, introducing deterministic truncation-like perturbations and stochastic perturbations in the acceptance probabilities. In (almost) all the cases we find a remarkable stability of the universal properties, even against large perturbations of the Markov Chain Monte Carlo algorithm, with critical exponents and scaling curves consistent with those of the standard XY model within statistical uncertainties.
@@ I Introduction
Since only few statistical models can be analytically solved, and mainly in two dimensions (see, e. g., Baxter-book ; IZ-book ; McCoy-book ), in statistical mechanics we often have to resort to numerical simulations. The Markov Chain Monte Carlo (MCMC) method is one of the most commonly used approaches to numerically investigate the equilibrium properties of statistical systems NB-book ; BH-book ; Berg-book ; LB-book ; Krauth-book ; DELL-book , since it provides stochastically exact estimates of average values whenever its application is not hindered by the presence of a sign problem LGSWSS-90 ; Troyer:2004ge ; Philipsen:2010gj ; Aarts:2015tyj .
@@
These theoretical results show that the MCMC approach is robust with respect to a large class of perturbations as far as we consider the first class of applications introduced above, however they seems to be of little use to understand whether universal properties are stable or not against perturbations of the MCMC algorithm. Since continuous phase transitions correspond to points of the phase diagram where the free energy is non-analytic in the thermodynamic limit, the closeness of the invariant probability distribution function of the perturbed MCMC to that of the exact MCMC, for fixed volume, is not enough to guarantee that the two Markov chains display the same critical behavior.
[…]
@@ II The XY model and the MCMC perturbations studied
[…]
@@
We consider two different update schemes: the single site Metropolis update and the single cluster update. In the single site Metropolis update we sweep through the lattice in lexicographic order, we generate a new random vector ${\bm{s}}^{(\mathrm{trial})}$, and accept the update ${\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}$ with probability
@@
where $\Delta E=H(\{{\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}\})-H(\{{\bm{s}}_{\bm{n}}\})$ is the energy change of the configuration induced by the substitution ${\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}$. In the single cluster update Wolff-89 we instead randomly select the direction ${\bm{v}}$ ($|{\bm{v}}|=1$) in the spins two dimensional space, and we denote by $R_{\bm{v}}({\bm{s}})$ the reflection with respect to plane orthogonal to ${\bm{v}}$:
@@
where $G^{\prime}_{\sigma}$ and $G_{\sigma}$ are two independent Gaussian random variables with zero average and variance $\sigma^{2}$. This perturbation can be taught of as a simple modeling of the case in which energies are not exactly computed but stochastically estimated. To mimic the effect of a finite machine precision we can instead use
@@ III.2 XY scaling curve
Table 3: Single site Metropolis update with Gaussian random energy perturbation (see Eq. (6)): estimates of the critical temperature for different values of $\sigma^{2}$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ).
@@
Table 4: Single site Metropolis update with energy truncation perturbation (see Eq. (7)): estimates of the critical temperature for different values of $\Lambda$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ).
In the lower panel of Fig. 1 we also show results related to the numerical estimate of $\mathcal{U}_{\omega}(R_{\xi})$, from which it can be seen that data obtained from the $L=8$ lattice systematically deviate from the other ones; this is why we included only $L\geq 16$ data in the fit. For what concerns the systematic errors of the $\mathcal{U}_{\omega}(R_{\xi})$ determination, we expect them to be larger than those affecting $\mathcal{U}(R_{\xi})$, but we have not data to quantify them in a robust way. For this reason, in the following analyses of the perturbed models, we will just use $\mathcal{U}(R_{\xi})$, and in particular the fact that $(U-\mathcal{U}(R_{\xi}))L^{\omega}$ converges to a scaling curve, without explicitly comparing this scaling curve with the estimate of $\mathcal{U}_{\omega}(R_{\xi})$ obtained from the XY model. Of course we could also directly compare data reported in the lower panel of Fig. 1 with the corresponding data of the perturbed model, but we would anyway have to fix the value of a non-universal parameter, see Sec. III.1. Since however statistical errors of the quantity $(U-\mathcal{U}(R_{\xi}))L^{\omega}$ will turn out to be quite large for the perturbed models, the result of this comparison (although positive) is not particularly significant, and will not be reported in the following.
@@ III.3.1 Single site Metropolis algorithm
To quantify the effect of the perturbation on non-universal quantities, for each value of $\sigma^{2}$ we estimate the critical temperature from the crossing points of $R_{\xi}$ and $U$ as a function of $\beta$. More precisely, we use the FSS reported in Eq. (14), with polynomial approximations of the functions $\mathcal{R}(x)$ and $\mathcal{R}_{\omega}(x)$, to fit data. In all the cases we fix $\nu$, $\omega$, $R_{\xi}^{*}$ and $U^{*}$ using the universal values of the XY universality class reported Tab. 1. Already the fact that the fit used to extract $\beta_{c}$ results in a $\chi^{2}/\mathrm{dof}\approx 1$ is thus a first indication that universal quantities are not affected by the stochastic noise. The residual dependence of the results on the degree of the polynomials used in the fit, or on the fit range adopted, has been added as a systematic error to the statistical error, and the final results are reported in Tab. 3. The good agreement between the estimates obtained using $R_{\xi}$ and those obtained using $U$ is an indication that systematical errors have been correctly taken into account.
@@ III.3.2 Single cluster update
Table 5: Single cluster update with Gaussian random energy perturbation of the cluster formation probability (see Eq. (8)): estimates of the critical temperature for different values of $\sigma^{2}$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ).
@@
Table 6: Single cluster update with truncation perturbation of the cluster formation probability (see Eq. (9)): estimates of the critical temperature for different values of $\Lambda$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ).
@@ IV Conclusions
For classical simulations performed on standard CPUs the topic investigated in this work is probably important mainly from the theoretical point of view, but it can be of more direct practical relevance in the context of quantum algorithms, or even to speed-up simulations using low-accuracy hardware.

### v2

@@ Abstract
We numerically investigate the stability of universal properties at continuous phase transitions against perturbations of the Markov Chain Monte Carlo algorithm used to simulate the system. We consider the three dimensional XY model as test bed, and both local (single site Metropolis) and global (single cluster) updates, introducing deterministic truncation-like perturbations and stochastic perturbations in the acceptance probabilities. In (almost) all the cases we find a remarkable stability of the universal properties, even against large perturbations of the Markov Chain Monte Carlo algorithm, with critical exponents and scaling curves consistent with those of the standard XY model within statistical uncertainties. Only for the single cluster update with very large truncation error does something different happen, but large scaling corrections prevent us from precisely assessing the critical properties of the transition, and, in particular, to understand whether the critical behavior observed corresponds to a known universality class.
@@ I Introduction
Since only few statistical models can be solved analytically, and mainly in two dimensions (see, e. g., Baxter-book ; IZ-book ; McCoy-book ), in statistical mechanics we often have to resort to numerical simulations. The Markov Chain Monte Carlo (MCMC) method is one of the most commonly used approaches to numerically investigate the equilibrium properties of statistical systems NB-book ; BH-book ; Berg-book ; LB-book ; Krauth-book ; DELL-book , since it provides stochastically exact estimates of average values whenever its application is not hindered by the presence of a sign problem LGSWSS-90 ; Troyer:2004ge ; Philipsen:2010gj ; Aarts:2015tyj .
@@
These theoretical results show that the MCMC approach is robust with respect to a large class of perturbations as long as we consider the first class of applications introduced above, however they seems to be of little use to understand whether universal properties are stable or not against perturbations of the MCMC algorithm. Since continuous phase transitions correspond to points of the phase diagram where the free energy is non-analytic in the thermodynamic limit, the closeness of the invariant probability distribution function of the perturbed MCMC to that of the exact MCMC, for fixed volume, is not enough to guarantee that the two Markov chains display the same critical behavior.
A simple example can be useful to clearly understand this point: if we consider the perturbation of the MCMC used to sample an Ising model obtained by introducing an external magnetic field $h$, it is immediate to see that, at fixed lattice volume, the probability distribution function of the $h\neq 0$ model is arbitrarily close to that of the $h=0$ model as far as we keep $|h|$ small enough. On the other hand, for whatever small but non-vanishing value of $|h|$ there is no phase transition if we keep $h$ fixed and perform the thermodynamic limit, although very large lattices may be needed to notice this fact in an actual numerical simulation.
@@ II The XY model and the MCMC perturbations studied
where the notation $\langle\phantom{n},\phantom{k}\rangle$ is used to specify that the sum extends on all neighboring sites of the lattice, and the partition function is thus given by ($\beta=1/(k_{B}T)$)
@@
We consider two different update schemes: the single site Metropolis update and the single cluster update. In the single site Metropolis update we sweep through the lattice in lexicographic order, we generate a new random vector ${\bm{s}}^{(\mathrm{trial})}$ by performing a random rotation of the spin ${\bm{s}}_{\bm{n}}$ (see later for more details), and accept the update ${\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}$ with probability
@@
where $\Delta E=H(\{{\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}\})-H(\{{\bm{s}}_{\bm{n}}\})$ is the energy change of the configuration induced by the substitution ${\bm{s}}_{\bm{n}}\to{\bm{s}}^{(\mathrm{trial})}$. To avoid the use of slow trigonometric functions the random rotation of the spin is performed by using the auxilliary functions
$c(\theta)=\frac{1}{\sqrt{1+\theta^{2}}}\ ,\quad s(\theta)=\theta c(\theta)\ ,$ (4)
with $\theta$ uniformly distributed in $(-2\pi,2\pi)$, and
$s^{\mathrm{trial}}_{x}=c(\theta)({\bm{s}}_{\bm{n}})_{x}-s(\theta)({\bm{s}}_{\bm{n}})_{y}\ ,$ (5)
$s^{\mathrm{trial}}_{y}=s(\theta)({\bm{s}}_{\bm{n}})_{x}+c(\theta)({\bm{s}}_{\bm{n}})_{y}\ .$
In the single cluster update Wolff-89 we instead randomly select the direction $\bm{v}$} ($|{\bm{v}}|=1$) in the spins two dimensional space, and we denote by $R_{\bm{v}}({\bm{s}})$ the reflection with respect to plane orthogonal to $\bm{v}$}:
@@
where $G^{\prime}_{\sigma}$ and $G_{\sigma}$ are two independent Gaussian random variables with zero average and variance $\sigma^{2}$. This perturbation can be thought of as a simple modeling of the case in which the energies of the initial and trial states are not exactly computed, but stochastically estimated. A different (not used) possibility would be to add to $\Delta E$ a single Gaussian random variable, to model the case in which directly $\Delta E$ is stochastically estimated. To mimic the effect of a finite machine precision we can instead use
@@ III.2 XY scaling curve
Table 3: Single site Metropolis update with Gaussian random energy perturbation (see Eq. (8)): estimates of the critical temperature for different values of $\sigma^{2}$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ). Note that the critical temperature is not universal, and there is no reason to expect $\beta_{c}$ not to vary with $\sigma^{2}$. Its variation is in fact an indication that the values of $\sigma^{2}$ used are large enough to have an effect on the system.
@@
Table 4: Single site Metropolis update with energy truncation perturbation (see Eq. (9)): estimates of the critical temperature for different values of $\Lambda$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ). Note that the critical temperature is not universal, and there is no reason to expect $\beta_{c}$ not to vary with $\Lambda$. Its variation is in fact an indication that the values of $\Lambda$ used are large enough to have an effect on the system.
In the lower panel of Fig. 1 we also show results related to the numerical estimate of $\mathcal{U}_{\omega}(R_{\xi})$, from which it can be seen that data obtained from the $L=8$ lattice in these high-accuracy simulations systematically deviate from the other ones; this is why we included only $L\geq 16$ data in the fit. For what concerns the systematic errors of the $\mathcal{U}_{\omega}(R_{\xi})$ determination, we expect them to be larger than those affecting $\mathcal{U}(R_{\xi})$, but we do not have data to quantify them in a robust way. For this reason, in the following analyses of the perturbed models, we will just use $\mathcal{U}(R_{\xi})$, and in particular the fact that $(U-\mathcal{U}(R_{\xi}))L^{\omega}$ converges to a scaling curve, without explicitly comparing this scaling curve with the estimate of $\mathcal{U}_{\omega}(R_{\xi})$ obtained from the XY model. Of course we could also directly compare data reported in the lower panel of Fig. 1 with the corresponding data of the perturbed model, but we would anyway have to fix the value of a non-universal parameter, see Sec. III.1. Since however statistical errors of the quantity $(U-\mathcal{U}(R_{\xi}))L^{\omega}$ will turn out to be quite large for the perturbed models, the result of this comparison (although positive) is not particularly significant, and will not be reported in the following.
@@ III.3.1 Single site Metropolis algorithm
To quantify the effect of the perturbation on non-universal quantities, for each value of $\sigma^{2}$ we estimate the critical temperature from the crossing points of $R_{\xi}$ and $U$ as a function of $\beta$. More precisely, we use the FSS reported in Eq. (16), with polynomial approximations of the functions $\mathcal{R}(x)$ and $\mathcal{R}_{\omega}(x)$, to fit data. In all the cases we fix $\nu$, $\omega$, $R_{\xi}^{*}$ and $U^{*}$ using the universal values of the XY universality class reported Tab. 1. Already the fact that the fit used to extract $\beta_{c}$ results in a $\chi^{2}/\mathrm{dof}\approx 1$ is thus a first indication that universal quantities are not affected by the stochastic noise. The residual dependence of the results on the degree of the polynomials used in the fit, on the fit range adopted, and on the minimum value of $L$ considered in the fit, has been added as a systematic error to the statistical error, and the final results are reported in Tab. 3. The good agreement between the estimates obtained using $R_{\xi}$ and those obtained using $U$ is an indication that systematic errors have been correctly taken into account.
@@ III.3.2 Single cluster update
Table 5: Single cluster update with Gaussian random energy perturbation of the cluster formation probability (see Eq. (10)): estimates of the critical temperature for different values of $\sigma^{2}$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ). Note that the critical temperature is not universal, and there is no reason to expect $\beta_{c}$ not to vary with $\sigma^{2}$. Its variation is in fact an indication that the values of $\sigma^{2}$ used are large enough to have an effect on the system.
@@
Table 6: Single cluster update with truncation perturbation of the cluster formation probability (see Eq. (11)): estimates of the critical temperature for different values of $\Lambda$, obtained by analyzing $R_{\xi}$ or $U$. For comparison the critical temperature of the XY model is 0.45416474(10)[7] (from Ref. Hasenbusch-19 ). Note that the critical temperature is not universal, and there is no reason to expect $\beta_{c}$ not to vary with $\Lambda$. Its variation is in fact an indication that the values of $\Lambda$ used are large enough to have an effect on the system.
@@ IV Conclusions
For classical simulations performed on standard CPUs the topic investigated in this work is probably important mainly from the theoretical point of view, but it can be of direct practical relevance in the context of quantum algorithms, and for simulations performed using low-accuracy hardware or surrogate energy functions.